5. Fill in the blank spaces in the table. Number of hours Total benefits Marginal benefits 100 178 210 playing baseball (in utils)(in utils) 2
Answer #1

Ans:

Number of hours playing baseball Total benefits ( in Utils ) Marginal benefits ( in Utils)
1 100 -
2 178 78
3 210 32

Explanation:

Marginal benefits = Change in Total benefit / Change in Number of hours playing baseball

= \Delta TB / \Delta Q

Marginal benefits from 2nd hour = ( 178-100 ) / ( 2-1) = 78 / 1 = 78

Marginal benefits from 3rd hour = ( 210 -178 ) / ( 3 -2 ) = 32 / 1 = 32